# Enhancing Clinical Trial Patient Matching through Knowledge Augmentation with Multi-Agents

### v1

@@ 4.3.1 Criterion-Level Performance
4.3.2 Trial-Level Performance
@@
The result of trail-levle analysis can be found in Table 2. At the trial level, where a patient’s eligibility is determined by satisfying all selected criteria, MAKA outperforms both Zero-shot and CoT. MAKA achieves the highest Accuracy (0.9306), Precision (0.6333)and F1-score (0.6552) but underperform CoT on Recall. The trail-level results further validate MAKA’s capability in maintaining a higher balance between identifying eligible patients and minimizing patient-filtering load
Our findings suggest that MAKA effectively can effectively identifies and addresses knowledge gaps in LLMs for trial-centric patient matching. By leveraging agent-augmented criteria, MAKA enhances LLMs’ understanding of clinical trial requirements. And it achieves a better performance on both minimizing manual selection efforts and accurately identifying eligible patients.
5 Limitations
@@
One significant limitation of this study lies in the dataset used for evaluation. The n2c2 dataset, the most commonly used, publicly available dataset for the cohort selection task, is relatively small in size and contains criteria that are neither particularly complex nor challenging for LLM-based processing. As a result, while MAKA is designed as a general framework capable of handling both simple and complex criteria, the dataset’s limitations restrict our ability to thoroughly evaluate its generalizability across a broader range of clinical trial requirements.
An additional observation from our experiments is that the Knowledge Probing Agent predominantly identified gaps in the criteria themselves—primarily issues of clarity or insufficient detail—rather than uncovering knowledge gaps in LLM. Moreover, the Navigation Agent consistently determined that all criteria could be sufficiently augmented by the Self-Augmentation Agent, with no need to employ external retrieval or online search augmentation. These findings suggest that the dataset’s simplicity may have limited the full utilization of MAKA’s modular capabilities.
6 Conclusion
In this paper, we introduced MAKA, a framework designed to automate decomposed patient-matching workflows by leveraging the advanced natural language understanding capabilities of LLMs. The framework’s modular design incorporates five key components—Knowledge Probing Agent, Navigation Agent, Knowledge Augmentation Agent, Supervision Agent, and Matching Agent—each contributing to systematic knowledge augmentation and robust decision-making in patient-trail matching settings.
We demonstrated that MAKA significantly improves patient-matching performance at both the criterion and trial levels. Compared to baseline approaches like Zero-shot and Chain of Thought (CoT), MAKA achieves notable gains in Accuracy, Precision, and F1-score, particularly when applied to complex or ambiguous criteria. By enriching clinical criteria with domain-specific insights, MAKA consistently outperforms alternatives while maintaining a balance between precision and recall. Further investigation of agent behaviors indicates that the performance improvements primarily stem from filling the knowledge gaps existing in the criteria, rather than in LLMs.
Our results highlight the practical benefits of augmenting LLM capabilities with a structured muti-agent workflow for knowledge enhancement. For example, MAKA-augmented criteria improves understanding of clinical trial requirements by aligning better with diverse patient note styles and heterogeneity of criterion. The results validate MAKA’s potential for real-world deployment in clinical trial workflows.
@@ 6 Conclusion
It is worth noting that the design of MAKA supports the future development of a human-AI collaboration system. For the trial-centric patient matching task, the volume of criterion-supervision performed by the supervisory agent is relatively small and can be efficiently handled by human experts with minimal time investment. Exploring how humans can further enhance and facilitate this system presents an interesting direction for future research.
@@
To address the limitations of this study and validate MAKA’s generalizability, future work could also focus on evaluating the framework using more challenging datasets, such as those involving oncology trial matching which contains highly complex criteria or those with a cohort of larger size. Testing MAKA in thess contexts would not only provide a more rigorous assessment of its capabilities but also demonstrate its potential to address critical real-world challenges in clinical trial recruitment.
Acknowledgments
@@
We gratefully acknowledge support from the OpenAI Research Access Program, which provided access to advanced language models that were instrumental in developing and evaluating the framework presented in this work. This support significantly facilitated the experiments and analyses conducted for this research.
@@ A.2 N2C2 Original Criteria
| CREATININE | Serum creatinine > upper limit of normal. |
A.3 Redefined Criteria
@@
A.4 MAKA-Augemented Criteria

### v2

@@ 4.3. Experiment Settings
4.4. Results of Experiment 1
4.4.1. Criterion-Level Performance:
Table 1 shows the performance of three approaches—Zero-shot,CoT, MAKAR—across multiple individual criteria. Our results indicate that while the CoT method slightly outperforms Zero-shot in Accuracy and Recall, MAKAR consistently achieves higher scores, particularly in Accuracy (0.909) and Precision(0.822) on average. The F1-score of MAKAR (0.828) suggests a balanced performance across criteria, although it falls slightly behind in a few specific cases such as ”ABDOMINAL” and ”English”.
The MAKAR-augmented criteria, presented in Table LABEL:tab:aug_criteria, demonstrate significant improvements over the baseline human-redefined criteria outlined in Table LABEL:table:redefined_criteria from Wornow et al. (2024). These enhancements primarily focus on providing explicit clarifications and detailed expansions of key concepts, enabling LLMs to better align with trial requirements. For example, MAKAR supplements normal ranges for criteria like ”HBA1C” and ”CREATININE” across diverse demographic groups. And they are also the criteria that benefit the most from MAKAR’s augmentation. Additionally, MAKAR provides a more comprehensive definition for criteria like ”MAJOR-DIABETES,” and use more widely recognized terminology, ”heart attack”, to explain clinical terms ”Myocardial Infarction (MI)”. These refinements improve the ability of LLMs to match patients from clinical notes written in a variety of styles, resulting in notable performance gains for these criteria.
4.4.2. Trial-Level Performance
@@
The result of trial-levle analysis can be found in Table 2. At the trial level, where a patient’s eligibility is determined by satisfying all selected criteria, MAKAR outperforms both Zero-shot and CoT. MAKAR achieves the highest Accuracy (0.9306), Precision (0.6333)and F1-score (0.6552) but underperform CoT on Recall. The trial-level results further validate MAKAR’s capability in maintaining a higher balance between identifying eligible patients and minimizing patient-filtering load
| Inclusion Criteria | |||||
| Metric | GPT-4o | o1 | o3-mini | CoT | MAKAR |
| Accuracy | 0.9000 | 0.9333 | 0.8533 | 0.9667 | 1.000 |
| Precision | 0.7884 | 0.8407 | 0.7211 | 0.9385 | 1.000 |
| Recall | 0.9020 | 0.9413 | 0.8151 | 0.9208 | 1.000 |
| F1-Score | 0.8281 | 0.8800 | 0.7516 | 0.9294 | 1.000 |
| Exclusion Criteria | |||||
| Metric | GPT-4o | o1 | o3-mini | CoT | MAKARCorr |
| Accuracy | 0.9000 | 0.9267 | 0.9133 | 0.9467 | 1.000 |
| Precision | 0.8693 | 0.8987 | 0.8840 | 0.9265 | 1.000 |
| Recall | 0.9090 | 0.9416 | 0.9253 | 0.9487 | 1.000 |
| F1-Score | 0.8845 | 0.9153 | 0.8999 | 0.9365 | 1.000 |
4.5. Results of Experiment 2
4.5.1. Results
Table 3 presents the evaluation results of our method on the in-house real-world dataset. We compared its performance against zero-shot prompting with GPT-4o, zero-shot prompting with two reasoning models (OpenAI-o3-mini and OpenAI-o1) and CoT strategy with GPT-4o. Our method achieves perfect scores across all four metrics.
Notably, for exclusion criteria, we report results for a corrected version of MAKAR (MAKARCorr). In the initial version, the model was not explicitly instructed on whether a given criterion was for inclusion or exclusion. Surprisingly, it was able to infer some exclusion criteria but focused on determining a patient’s overall eligibility for a trial rather than assessing whether a specific criterion was met. For instance, if a patient was 75 years old, they should have been marked as ”MET” for an exclusion criterion of ”greater than 70 years old”. While MAKAR correctly identified the age range and provided a logical reasoning process, it mistakenly output ”Not MET”—interpreting the task as determining overall trial eligibility rather than verifying the criterion itself. The corrected version addresses this by explicitly instructing the model to evaluate each criterion independently rather than making a broader eligibility judgment.
MAKAR outperforms all the benchmarks and the CoT follows closely, achieving 96.67% accuracy and 92.94% F1-score for inclusion criteria and 94.67% accuracy with 93.65% F1-score for exclusion criteria. o3-mini, a smaller model, is outperformed by o1 but shows competitive performance on exclusion criteria with zero-shot method with GPT-4o. Interestingly, Experiment 2 yields better overall performance than Experiment 1, despite its more knowledge-intensive criteria. We hypothesize that this is due to differences in the labeling process—the n2c2 challenge involves more subjective judgment, while our setup explicitly defines insufficient information as ”NOT MET”.
4.5.2. Error Analysis
Despite MAKAR achieving 100% accuracy in the second experiment, we manually checked the reasoning results of all 300 patient-criterion pair and identified two reasoning errors that did not affect the final output.
First, in Exclusion Criterion 13 (Patients with ongoing indications for cardioprotective medication—ACE inhibitors, ARBs, and/or beta-blockers), MAKAR reasoned: ”The patient is on metoprolol, a beta-blocker, suggesting some ongoing indication for its use. However, without explicit information on the specific condition necessitating this medication (e.g., arrhythmia, prophylactic use post-cardiac event other than MI, or another reason), the justification remains uncertain.” This reasoning was overly conservative, but the Supervision Matching Agent correctly classified it as MET, aligning with expert judgment.
Second, in Inclusion Criterion 5 (Histologically confirmed breast cancer with an invasive component 20 mm and/or morphologically confirmed spread to regional lymph nodes(stage cT2-cT4 with any cN, or cN1-cN3 with any cT).), MAKAR reasoned: ”The documentation mentions a suspicious breast nodule at the right breast (1.4 × 1.5 × 0.5 cm), which measures more than 20 mm in one dimension.” This was an obvious error, as the maximum dimension stated is only 1.5 cm. The mistake may come from over-reasoning, possibly computing the largest possible diagonal as . However, the final decision remained correct since the criterion lymph node involvement still holds.
4.6. Ablation
In this section, we validate the contribution of each module in our framework. First, we remove the knowledge augmentation module to assess the reasoning module (MAKA-reason) using the original criteria, which essentially mirrors the CoT approach in Table 3. Next, we remove the reasoning module and evaluate the knowledge augmentation module (MAKA-aug) independently by performing zero-shot matching with the augmented criteria. The results, presented in Table 4, indicate that removing the reasoning module has a greater negative impact on inclusion criteria matching (F1-score: 88.37% vs. 92.94%), while removing the augmentation module has a larger effect on exclusion criteria (F1-score: 93.65% vs. 94.41%).
Although we did not involve any reasoning-LLM-based agents, our approach still shows impressive reasoning capabilities. This can be attributed to the well-structured augmented criteria and the CoT strategy employed by the reasoning agent. For instance, when evaluating whether a patient meets the Inclusion Criterion 14 (Postmenopausal hormone receptor-positive patients), MAKA reasoned: ”…the record does not explicitly state whether the breast cancer was hormone receptor positive. However, the use of tamoxifen, which is specifically prescribed for estrogen receptor-positive breast cancer by blocking estrogen receptors, strongly implies that her breast cancer was hormone receptor positive.” Both MAKA and MAKA-reason successfully made the correct match, while MAKA-aug failed to capture this implicit relationship.
Additionally, we also observed cases where knowledge augmentation improved performance. For example, in Exclusion Criterion 7, it mentions: ”…all histological lesions were HER2 1+ or 2+, not detected by FISH amplification, and HR positive according to ASCO guidelines”. MAKA-aug expanded the definitions from American Society of Clinical Oncology(ASCO) guidelines, made correct matches for all 10 patients. In contrast, MAKA-reason correctly identified only 8 out of 10.
| Inclusion Criteria | |||
| Metric | MAKAR-reason | MAKAR-aug | MAKAR |
| Accuracy | 0.9667 | 0.9333 | 1.000 |
| Precision | 0.9385 | 0.8387 | 1.000 |
| Recall | 0.9208 | 0.9612 | 1.000 |
| F1-Score | 0.9294 | 0.8837 | 1.000 |
| Exclusion Criteria | |||
| Metric | MAKAR-reason | MAKAR-aug | MAKARCorr |
| Accuracy | 0.9467 | 0.9533 | 1.000 |
| Precision | 0.9265 | 0.9360 | 1.000 |
| Recall | 0.9487 | 0.9534 | 1.000 |
| F1-Score | 0.9365 | 0.9441 | 1.000 |
5. Discussion
One interesting observation from our experiments is that the Router Agent consistently determined that all criteria could be augmented by the Self-Augmentation Agent, without requiring external retrieval or online search augmentation. This was true even for breast cancer-related criteria, which are typically complex and challenging for individuals without professional training.
@@ 5. Discussion
This relates to a limitation of our study, which comes from the dataset used for evaluation. Both the n2c2 dataset and the real-world dataset are relatively small and contain criteria that are not particularly challenging for LLM-based processing, allowing all benchmark methods to achieve strong performance. While MAKAR is designed as a general framework capable of handling both simple and complex criteria, these dataset constraints limit our ability to explore the knowledge boundaries of LLMs and fully assess the framework’s generalizability across a wider range of clinical trial requirements.
Moreover, this observation raises a discussion about the specific improvements our framework brings to baseline LLMs. While the framework is designed to bridge the knowledge gap between trial criteria and LLMs’ existing knowledge, our experiments indicate that LLMs can already expand and interpret the criteria in our datasets accurately. Instead, the primary benefit of knowledge augmentation lies in explicitly organizing relevant concepts and definitions in a well-structured format. By clearly outlining these key elements, the augmentation provides a type of breakdown of the subsequent reasoning steps, leading to improved overall performance.
However, despite the dataset limitations, the effectiveness of the augmentation module remains significant. One of its key strengths is its potential to interface with diverse data sources, particularly those that are not publicly available, such as paid subscription-based medical databases or proprietary clinical trial repositories. Furthermore, it plays a crucial role in integrating the most up-to-date medical knowledge, including newly developed treatments and evolving FDA guidelines. This capability ensures that the framework remains adaptable and relevant, even when dealing with rapidly changing medical landscapes where LLMs alone may lack the latest information. By systematically incorporating external knowledge, the augmentation module enhances the contextual relevance of trial matching, making it particularly valuable in real-world applications where timely and comprehensive knowledge retrieval is essential.
6. Conclusion
In this study, we introduced MAKAR, a scalable multi-agent framework designed to enhance patient-trial matching by combining knowledge augmentation with structured reasoning. By integrating an automated augmentation module with a reasoning module, MAKAR refines complex trial criteria and systematically evaluates patient eligibility. Our experimental results demonstrate that MAKAR outperforms existing benchmarks, including the state-of-the-art closed-source reasoning LLMs, and achieve 100% accuracy in a real-world offline test. And it also achieves up to an 8% improvement on complex criteria in a public dataset. Beyond performance gains, our findings emphasize the importance of explicitly structuring and expanding trial criteria to bridge the knowledge gap in AI-driven clinical reasoning.
@@ 6. Conclusion
It is worth noting that the design of MAKAR supports the future development of a human-AI collaboration system. For the trial-centric patient matching task, the workload of criterion-supervision performed by the critic agent is one-time per trial and relatively small and can be efficiently handled by human experts with minimal time investment. Exploring how humans can further enhance and facilitate this system presents an interesting direction for future research.
@@
To address the limitations of this study and validate MAKAR’s generalizability, future work could also focus on evaluating the framework using more challenging datasets, such as those involving newly launched oncology trial that required up-to-date medical knowledge. Testing MAKAR in thess contexts would not only provide a more rigorous assessment of its capabilities but also demonstrate its potential to address critical real-world challenges in clinical trial recruitment.
7. Acknowledgments
@@
We gratefully acknowledge support from the OpenAI Research Access Program, which provided access to advanced language models that were instrumental in developing and evaluating the framework presented in this work.
@@ A.2. N2C2 Original Criteria
| CREATININE | Serum creatinine ¿ upper limit of normal. |
A.3. Redefined N2C2 Criteria
@@
A.4. MAKAR-Augemented N2C2 Criteria
@@
A.5. Sampled ClinicalTrial.gov Criteria
| Inclusion Criteria | Exclusion Criteria |
| 1. Locally advanced or metastatic breast cancer confirmed by histopathology. | 1. Patients known or suspected intolerance or hypersensitivity to main ingredient or any of the excipients of SNB-101. |
| 2. Received 2 previous lines of anti-tumor treatment and developed resistance to standard treatment. | 2. Patients with bilateral invasive breast cancers. |
| 3. Patients with locally advanced or metastatic advanced solid tumors confirmed by histology or cytology (including but not limited to triple-negative breast cancer, gastric cancer, colorectal cancer); Patients with 1 line of standard treatment failure (disease progression after treatment or intolerability of toxic side effects of treatment), or no standard treatment, or unable to receive standard treatment. | 3. Age 18 years old or 70 years old. |
| 4. At least 1 measurable lesion per Response Evaluation Criteria in Solid Tumors version 1.1 and previously treated lesions with radiotherapy or focal therapy and no progression cannot be included as target lesion for assessment. | 4. Patients with standard metallic contraindications to CMR or estimated glomerular filtration rate 30 mL/min/1.73 m². |
| 5. Histologically confirmed breast cancer with an invasive component measuring 20 mm and/or with morphologically confirmed spread to regional lymph nodes (stage cT2-cT4 with any cN, or cN1-cN3 with any cT). | 5. History of hypersensitivity or contraindication to TTF. |
| 6. Known HER2-positive breast cancer defined as an IHC status of 3+. If IHC is 2+, a positive in situ hybridization (FISH, CISH, or SISH) test is required by local laboratory testing. | 6. Implanted pacemaker, defibrillator, or other electrical medical devices. |
| 7. Patients with breast cancer planned to receive Anthracycline and Cyclophosphamide chemotherapy. | 7. Bilateral breast cancer (including multifocal breast cancer. All histological lesions were HER2 1+ or 2+, not detected by FISH amplification and HR positive according to ASCO guidelines). |
| 8. Patients treated for hormone-dependent localized breast cancer requiring adjuvant hormonal therapy (HT). | 8. Received any form of anti-tumor therapy (chemotherapy, radiotherapy, molecular targeted therapy, endocrine therapy, etc.). |
| 9. Patients treated with tamoxifen for a maximum of 1 to 3 years. | 9. Breast cancer without histopathological diagnosis. |
| 10. Histologically confirmed ER+/HER2- early-stage resected invasive breast cancer at high or intermediate risk of recurrence, , based on clinical-pathological risk features, as defined in the protocol. | 10. Known personal history of ductal carcinoma in situ (DCIS) or invasive breast cancer. |
| 11. Completed adequate (definitive) locoregional therapy (surgery with or without radiotherapy) for the primary breast tumour(s), with or without (neo)adjuvant chemotherapy. | 11. Prior systemic treatment for any malignancy. |
| 12. Completed at least 2 years but no more than 5 years (+3 months) of adjuvant ET (+/- CDK4/6 inhibitor). | 12. Significant medical comorbidities as per investigator evaluation. |
| 13. Patients receiving letrozole for more than two months. | 13. Patients with ongoing indications for the cardioprotective medication - ACE inhibitors, ARBs and/or beta-blockers |
| 14. Postmenopausal hormone receptor-positive patients. | 14. Intermediate or high-grade ductal carcinoma in situ. |
| 15. Patients with N0 disease or those who have undergone targeted axillary detection with a good response. | 15. Invasive carcinoma. |